OpenStudy (anonymous):

Ill give medal, fan testimonial, and virtual hug for help. Note: Remember to show all of the steps that you use to solve the problem. You can use the comments field to explain your work. Your teacher will review each step of your response to ensure you receive proper credit for your answer. Suppose a parabola has vertex (–4, 7) and also passes through the point (–3, 8). Write the equation of the parabola in vertex form.

OpenStudy (anonymous):

looks like @hartnn is on the job, ur def get an answer, no point in helping u

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

my life is over :,(

OpenStudy (anonymous):

y=a(x−h)2+k That's the formula. Enter the points and solve for a: 8=a((−3)+4)2+7 8=a(1)2+7 1 = a Therefore the vertex form of the parabola is: y=(x+4)2+7

OpenStudy (anonymous):

yeah sorry i do not get that at all

OpenStudy (anonymous):

srry... in the first line, its y=a (x - h)to the second power + k